# HopCast: Calibration of Autoregressive Dynamics Models

### v4

@@ Abstract
Deep learning models are often trained to approximate dynamical systems that can be modeled using differential equations. Many of these models are optimized to predict one step ahead; such approaches produce calibrated one-step predictions if the predictive model can quantify uncertainty, such as Deep Ensembles. At inference time, multi-step predictions are generated via autoregression, which needs a sound uncertainty propagation method to produce calibrated multi-step predictions. This work introduces an alternative Predictor-Corrector approach named HopCast that uses Modern Hopfield Networks (MHN) to learn the errors of a deterministic Predictor that approximates the dynamical system. The Corrector predicts a set of errors for the Predictor’s output based on a context state at any timestep during autoregression. The set of errors creates sharper and well-calibrated prediction intervals with higher predictive accuracy compared to baselines without uncertainty propagation. The calibration and prediction performances are evaluated across a set of dynamical systems. This work is also the first to benchmark existing uncertainty propagation methods based on calibration errors.
@@ 1 Introduction
Approximating dynamical systems that can be modeled using differential equations has many applications in systems biology (Wang et al., 2021; Brauer & Kribs, 2016), control (Moerland et al., 2023; Lu et al., 2021), economics (Tu, 2012), and cyber-physical systems (Bilal Shahid & Fleming, 2025; Shahid et al., 2024; Robison et al., 2024). Many deep learning approaches to approximate dynamical systems optimize the model to predict one step ahead during training (Khansari-Zadeh & Billard, 2011; Coates et al., 2008; Ko et al., 2007). The multi-step predictions are generated via autoregression at inference time (Lu et al., 2021). In doing so, the model error accumulates over time, resulting in multi-step predictions diverging from the ground truth (Venkatraman et al., 2015; Janner et al., 2021). Hence, it is imperative to develop methods to generate accurate and calibrated multi-step predictions.
State-of-the-art methods for generating accurate and calibrated predictions include Probabilistic Ensembles (Deep Ensembles) (Lakshminarayanan et al., 2017). To approximate the dynamics typically associated with differential equations, these models are optimized to predict one step ahead during training and then autoregressively generate multi-step predictions at inference time. To generate calibrated multi-step predictions, Probabilistic Ensembles need sound uncertainty propagation methods such as Trajectory Sampling (Chua et al., 2018).
In contrast, we introduce a Predictor-Corrector mechanism to produce accurate and calibrated multi-step predictions for dynamical systems without the need for uncertainty propagation. The Predictor is optimized to predict a one-step ahead point estimate of the next state of the system during training and produce multi-step predictions via autoregression at inference time. Our Corrector is based on the idea that similar states of a system lead to similar errors. This idea was leveraged by Auer et al. (2023) to build calibrated prediction intervals for one-step prediction utilizing Modern Hopfield Networks (MHN). However, the same state may not lead to the same error for multi-step predictions for dynamical systems. In such situations, the error also depends on how the system arrives at that state, i.e., where the system started and when it reached the state. To that end, we introduce the concept of context state, assuming that similar context states lead to similar errors. The Corrector, utilizing MHN, learns similarity between context states in terms of errors. At any timestep during autoregression, the Predictor predicts, and the Corrector generates prediction intervals for the prediction. To control the width of prediction intervals for calibration, we introduce the concept of Attention Span. We name our approach HopCast.
@@
  • lower calibration and prediction error across several benchmarks, without the use of complex uncertainty propagation techniques. Unlike uncertainty propagation methods, HopCast creates sharper prediction intervals based on context state instead of accumulating uncertainty over time during autoregression.
@@ 2 Related Work
Bayesian Methods. In Bayesian models, a distribution over the parameters of the underlying model is learned, called posterior distribution. At inference time, multiple samples of parameters can be taken from the posterior distribution using an accurate sampling method, such as Hamiltonian Monte Carlo (HMC) (Betancourt, 2017). These samples serve as multiple predictive models and give us the desired diversity in the predictive variable of interest. A notable method from this domain is Bayes by Backprop (Blundell et al., 2015), which learns an approximate posterior distribution over the weights of a feedforward model via Variational Inference (VI) by optimizing the evidence lower bound (ELBO) (Kingma et al., 2015).
@@
Ensemble Methods. In classical ensemble methods, an ensemble of models is trained with variations in the data for each model in the ensemble to train robust predictors (Davison & Hinkley, 1997). Deep Ensembles, however, are constructed using modern deep learning models with random initializations in the parameter space. These over-parametrized models, when initialized randomly, converge to different local minima, giving us the desired diversity in the predictive variable (Fort et al., 2019) These models were first presented by Lakshminarayanan et al. (2017) in the context of predictive uncertainty estimation. A notable extension of that is Neural Ensemble Search (Zaidi et al., 2021), which uses different architectures for models in an ensemble and improves upon deep ensembles. The Deep Ensemble is a scalable way of quantifying uncertainty compared to Bayesian methods (Wilson & Izmailov, 2020). Hence, these are used as baseline models with different uncertainty propagation methods in the current work.
@@
Model Calibration. Calibration has been studied widely for classification models. In binary classification, Platt Scaling (Platt et al., 1999) and isotonic regression (Niculescu-Mizil & Caruana, 2005) have been used successfully for recalibration. There are extensions of such works to multi-class classification problems (Zadrozny & Elkan, 2002). In the context of regression, Gneiting & Raftery (2007) proposed several proper scoring rules to evaluate the calibration of a probabilistic model for continuous variables. Those scoring rules have been used in the literature as loss functions, for example, continuous ranked probability score (CRPS) (Gasthaus et al., 2019). Calibration has also been discussed in the literature on probabilistic forecasting, mainly in the context of meteorology (Gneiting & Raftery, 2005), resulting in specialized calibration systems (Raftery et al., 2005). An approach called calibrated regression was proposed by Kuleshov et al. (2018), which used isotonic regression to recalibrate Bayesian models. In contrast, our work turns a deterministic Predictor into a calibrated Predictor via a Predictor-Corrector mechanism. Also, we focus on Predictors that autoregressively generate the next state of a dynamical system, a setting rarely addressed in the model calibration literature to the best of our knowledge.
@@ 3 Problem Description
Consider a dynamical system that can be defined using differential equations, such as Lotka-Volterra (LV). The LV system is used throughout the following discussion for illustrative purposes without loss of generality; subsequent sections show results for a variety of systems. LV dynamics can be written as:
@@
The state of this system can be defined as: $\mathbf{\tilde{s}}=(x,y)$. We assume a set of ground-truth $N$ trajectories of $T$ timesteps $I_{N}=\{\tau_{n}\}_{n=1}^{N}$ where $\tau_{n}=\{(\mathbf{s}_{t}^{n},\mathbf{s}_{t+1}^{n})_{t}\}_{t=1}^{T}$ and $\mathbf{s}_{t}^{n}=(x_{t}^{n},y_{t}^{n})$. This dataset can be generated by solving LV with a numerical ODE solver such as $\mathtt{solve\_ivp}$ from $\mathtt{scipy}$ with random initial conditions. To simulate measurement noise, we perturb each state post-integration with additive Gaussian noise, i.e., $\mathbf{s}=\mathbf{\tilde{s}}+\mathbf{\epsilon}$, where $\mathbf{\epsilon}\sim\mathcal{N}(0,\sigma*\Sigma)$ and $\Sigma=\text{diag}(\sigma_{x},\sigma_{y})$. The $\sigma$ is the scaling factor. We name $I_{N}$ an Integrated dataset. We assume that there is a Predictor $B$ (trained on $I_{N}$) that predicts $\mathbf{\bar{s}}_{t+1}$  [^1] The overhead bar over a variable shows that it belongs to Autoregressive dataset. given $\mathbf{\bar{s}}_{t}$, autoregressively generating a predicted set of $N$ trajectories starting from the same initial conditions. Precisely, $A_{N}=\{\bar{\tau}_{n}\}_{n=1}^{N}$ where $\bar{\tau}_{n}=\{(\mathbf{\bar{s}}_{t}^{n},\mathbf{\bar{s}}_{t+1}^{n})_{t}\}_{t=1}^{T}$ and $\mathbf{\bar{s}}_{t}^{n}=(\bar{x}_{t}^{n},\bar{y}_{t}^{n})$. We call $A_{N}$ an Autoregressive dataset. The error ($\mathbf{e}_{t}^{n}$) is the difference between $t^{th}$ timestep of $n^{th}$ trajectory from Integrated and Autoregressive, that is, $\mathbf{e}_{t}^{n}=(e_{x_{t}^{n}},e_{y_{t}^{n}})$ where $\mathbf{e}_{t}^{n}=\mathbf{s}_{t}^{n}-\bar{\mathbf{s}}_{t}^{n}$. Therefore, the set of errors of $N$ trajectories is $F_{N}=\{\{\mathbf{e}_{t}^{n}\}_{t=1}^{T}\}_{n=1}^{N}$.
@@ 4 HopCast
The proposed methodology, HopCast, consists of a Predictor-Corrector mechanism shown in Fig. 1. At any timestep ($t$) during autoregression, the Predictor $B$ predicts the next state of the system $(\bar{x}_{t+1},\bar{y}_{t+1})$ given previous state $(\bar{x}_{t},\bar{y}_{t})$. The Corrector consists of correction models ($M_{x}$ and $M_{y}$), one for each output of the Predictor $B$. Particularly, each correction model ($M$) consists of an Encoder ($m$) and an MHN (Ramsauer et al., 2020; Auer et al., 2023). The Encoder ($m$) is a fully connected feedforward model. Architectural details are included in Appendix D.1. We frame correction as a pattern retrieval task (Ramsauer et al., 2020). Precisely, $M_{x}$ and $M_{y}$ take in the context state $(\bar{x}_{1},\bar{y}_{1},\bar{x}_{t},\bar{y}_{t},t)$ and outputs $E_{\bar{x}}$ and $E_{\bar{y}}$, respectively. $E_{\bar{x}}$ and $E_{\bar{y}}$ contains errors of similar context states from past. The context state is the input of Predictor $B$ $(\bar{x}_{t},\bar{y}_{t})$ augmented with the initial condition $(\bar{x}_{1},\bar{y_{1}})$ and time step ($t$). The $E_{\bar{x}}$ & $E_{\bar{y}}$ are used to derive prediction intervals.
@@
The rest of this section has three sub-sections. Section 4.1 discusses the shuffling of $I_{N}$, $A_{N}$, & $F_{N}$ datasets to train correction models for pattern retrieval tasks. Section 4.2 then discusses the training of the correction model ($M_{x}$), and section 4.3 describes the inference with $M_{x}$ where we generate the set of errors ($E_{\bar{x}}$) given the context state.
@@ 4.1 Shuffled Data
Shuffled Queries ($\mathcal{S}_{Q}$). Before training the correction model ($M_{x}$), the states of the system from Autoregressive ($A_{N}$) dataset are augmented with context information, that is, initial condition and time step. Appendix D.2 emphasizes the importance of adding initial condition as a context. Specifically, the $t^{th}$ timestep of $n^{th}$ trajectory $\bar{\tau}_{n}$ from $A_{N}$, i.e., $(\bar{x}_{t}^{n},\bar{y}_{t}^{n})$, will be $\bar{c}_{t}^{n}=(\bar{x}_{1}^{n},\bar{y}_{1}^{n},\bar{x}_{t}^{n},\bar{y}_{t}^{n},t)$ after adding context information. The tuple $(\bar{x}_{1}^{n},\bar{y}_{1}^{n})$ is the initial condition of trajectory $\bar{\tau}_{n}$. The $\bar{c}_{t}^{n}$ denotes the context state associated with the $t^{th}$ timestep of $n^{th}$ trajectory. The set of all context states constructed from dataset $A_{N}$ is $\{\{\bar{c}_{t}^{n}\}_{t=1}^{T}\}_{n=1}^{N}$. This nested set of context states is flattened and shuffled randomly to form a new set $\mathcal{S}_{Q}=\text{Shuffle}\left(\bigcup_{n=1}^{N}\{\bar{c}_{t}^{n}\}_{t=1}^{T}\right)$. Appendix D.5 highlights the importance of shuffling the data.
Shuffled Keys ($\mathcal{S}_{K}$). Likewise, we construct the context states $\{\{c_{t}^{n}\}_{t=1}^{T}\}_{n=1}^{N}$ for Integrated ($I_{N}$) dataset as well, where $c_{t}^{n}=(x_{1}^{n},y_{1}^{n},x_{t}^{n},y_{t}^{n},t)$. Similar to $\mathcal{S}_{Q}$, we flatten the nested set $\{\{c_{t}^{n}\}_{t=1}^{T}\}_{n=1}^{N}$ and randomly shuffle it to form $\mathcal{S}_{K}=\text{Shuffle}\left(\bigcup_{n=1}^{N}\{c_{t}^{n}\}_{t=1}^{T}\right)$.
@@
Once we have $\mathcal{S}_{Q}$, $\mathcal{S}_{K}$, $\mathcal{S}_{V_{x}}$, $\mathcal{S}_{V_{y}}$, they are split into $\mathcal{S}_{Q}^{\text{Train}}$, $\mathcal{S}_{K}^{\text{Train}}$, $\mathcal{S}_{V_{x}}^{\text{Train}}$, $\mathcal{S}_{V_{y}}^{\text{Train}}$ and $\mathcal{S}_{Q}^{\text{Inference}}$, $\mathcal{S}_{K}^{\text{Inference}}$, $\mathcal{S}_{V_{x}}^{\text{Inference}}$, $\mathcal{S}_{V_{y}}^{\text{Inference}}$ with an 80/20 split. To train $M_{x}$, we need dataset $\mathcal{D}_{x}=\{(\mathbf{Q},\mathbf{K},\mathbf{V}_{x})_{i}\}$ comprising triplets $(\mathbf{Q},\mathbf{K},\mathbf{V}_{x})$. Each triplet of $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}_{x}$ are constructed by sampling $\mathtt{S_{L}}$ (Sequence Length) number of elements from sets $\mathcal{S}_{Q}^{\text{Train}}$, $\mathcal{S}_{K}^{\text{Train}}$, and $\mathcal{S}_{V_{x}}^{\text{Train}}$, respectively. For instance, one triplet with $\mathtt{S_{L}}=5$ is: $\mathbf{Q}=(\bar{c}_{4}^{3},\bar{c}_{5}^{1},\bar{c}_{7}^{5},\bar{c}_{2}^{3},\bar{c}_{1}^{7})^{\top}$, $\mathbf{K}=(c_{4}^{3},c_{5}^{1},c_{7}^{5},c_{2}^{3},c_{1}^{7})^{\top}$, and $\mathbf{V}_{x}=(e_{x_{4}^{3}},e_{x_{5}^{1}},e_{x_{7}^{5}},e_{x_{2}^{3}},e_{x_{1}^{7}})^{\top}$. The error $e_{x_{4}^{3}}=x_{4}^{3}-\bar{x}_{4}^{3}$ denotes the error of Predictor $B$ at $4^{th}$ timestep of $3^{rd}$ trajectory. Likewise, we construct $\mathcal{D}_{y}=\{(\mathbf{Q},\mathbf{K},\mathbf{V}_{y})_{i}\}$ using $\mathcal{S}_{Q}^{\text{Train}}$, $\mathcal{S}_{K}^{\text{Train}}$ and $\mathcal{S}_{V_{y}}^{\text{Train}}$. The 20% split with inference data is used for final evaluation.

### v5

@@ Abstract
Deep learning models are often trained to approximate dynamical systems that can be modeled using differential equations. Many of these models are optimized to predict one-step ahead; such approaches produce calibrated one-step predictions if the predictive model can quantify uncertainty, such as Deep Ensembles. At inference time, multi-step predictions are generated via autoregression, which needs a sound uncertainty propagation method to produce calibrated multi-step predictions. This work introduces an alternative Predictor-Corrector approach named HopCast that uses Modern Hopfield Networks (MHN) to learn the errors of a deterministic Predictor that approximates the dynamical system. The Corrector predicts a set of errors for the Predictor’s output based on a context state at any timestep during autoregression. The set of errors creates sharper and well-calibrated prediction intervals with higher predictive accuracy compared to baselines without uncertainty propagation. The calibration and prediction performances are evaluated across a set of dynamical systems. This work is also the first to benchmark existing uncertainty propagation methods based on calibration errors. We also evaluate HopCast as a substitute for Deep Ensembles within a model-based reinforcement learning planner, demonstrating improved performance across multiple control tasks.
@@ 1 Introduction
Approximating dynamical systems that can be modeled using differential equations has many applications in systems biology (wang2021applications; brauer2016dynamical), control (moerland2023model; lu2021revisiting), economics (tu2012dynamical), and cyber-physical systems (bilal2025toward; shahid2024uncertainty; robison2024differential). Many deep learning approaches to approximate dynamical systems optimize the model to predict one-step ahead during training (khansari2011learning; coates2008learning; ko2007gp). The multi-step predictions are generated via autoregression at inference time (lu2021revisiting). In doing so, the model error accumulates over time, resulting in multi-step predictions diverging from the ground truth (venkatraman2015improving; janner2021offline). Hence, it is imperative to develop methods to generate accurate and calibrated multi-step predictions.
State-of-the-art methods for generating accurate and calibrated predictions include Probabilistic Ensembles (Deep Ensembles) (lakshminarayanan2017simple). To approximate the dynamics typically associated with differential equations, these models are optimized to predict one-step ahead during training and then autoregressively generate multi-step predictions at inference time. To generate calibrated multi-step predictions, Probabilistic Ensembles need sound uncertainty propagation methods such as Trajectory Sampling (chua2018deep).
In contrast, we introduce a Predictor-Corrector mechanism to produce accurate and calibrated multi-step predictions for dynamical systems without the need for uncertainty propagation. The Predictor is optimized to predict a one-step ahead point estimate of the next state of the system during training and produce multi-step predictions via autoregression at inference time. Our Corrector is based on the idea that similar states of a system lead to similar errors. This idea was leveraged by auer2023conformal to build calibrated prediction intervals for one-step prediction utilizing Modern Hopfield Networks (MHN). However, the same state may not lead to the same error for multi-step predictions for dynamical systems. In such situations, the error also depends on how the system arrives at that state, i.e., where the system started and when it reached the state. To that end, we introduce the concept of context state, assuming that similar context states lead to similar errors. The Corrector, utilizing MHN, learns similarity between context states in terms of errors. At any timestep during autoregression, the Predictor predicts, and the Corrector generates prediction intervals for the prediction. To control the width of prediction intervals for calibration, we introduce the concept of Attention Span. We name our approach HopCast.
@@
  • lower calibration and prediction error across several benchmarks, without the use of complex uncertainty propagation techniques. Unlike uncertainty propagation methods, HopCast creates sharper prediction intervals based on context state instead of accumulating uncertainty over time during autoregression. We also test HopCast as a replacement for Probabilistic Ensembles in a model-based reinforcement learning planner, achieving strong performance on standard control tasks.
@@ 2 Related Work
Bayesian Methods. In Bayesian models, a distribution over the parameters of the underlying model is learned, called posterior distribution. At inference time, multiple samples of parameters can be taken from the posterior distribution using an accurate sampling method, such as Hamiltonian Monte Carlo (HMC) (betancourt2017conceptual). These samples serve as multiple predictive models and give us the desired diversity in the predictive variable of interest. A notable method from this domain is Bayes by Backprop (blundell2015weight), which learns an approximate posterior distribution over the weights of a feedforward model via Variational Inference (VI) by optimizing the evidence lower bound (ELBO) (kingma2015variational).
@@
Ensemble Methods. In classical ensemble methods, an ensemble of models is trained with variations in the data for each model in the ensemble to train robust predictors (davison1997bootstrap). Deep Ensembles, however, are constructed using modern deep learning models with random initializations in the parameter space. These over-parametrized models, when initialized randomly, converge to different local minima, giving us the desired diversity in the predictive variable (fort2019deep) These models were first presented by lakshminarayanan2017simple in the context of predictive uncertainty estimation. A notable extension of that is Neural Ensemble Search (zaidi2021neural), which uses different architectures for models in an ensemble and improves upon deep ensembles. The Deep Ensemble is a scalable way of quantifying uncertainty compared to Bayesian methods (wilson2020bayesian). Hence, these are used as baseline models with different uncertainty propagation methods in the current work.
@@
Model Calibration. Calibration has been studied widely for classification models. In binary classification, Platt Scaling (platt1999probabilistic) and isotonic regression (niculescu2005predicting) have been used successfully for recalibration. There are extensions of such works to multi-class classification problems (zadrozny2002transforming). In the context of regression, gneiting2007strictly proposed several proper scoring rules to evaluate the calibration of a probabilistic model for continuous variables. Those scoring rules have been used in the literature as loss functions, for example, continuous ranked probability score (CRPS) (gasthaus2019probabilistic). Calibration has also been discussed in the literature on probabilistic forecasting, mainly in the context of meteorology (gneiting2005weather), resulting in specialized calibration systems (Raftery2005BMA). An approach called calibrated regression was proposed by kuleshov2018accurate, which used isotonic regression to recalibrate Bayesian models. In contrast, our work turns a deterministic Predictor into a calibrated Predictor via a Predictor-Corrector mechanism. Also, we focus on Predictors that autoregressively generate the next state of a dynamical system, a setting rarely addressed in the model calibration literature to the best of our knowledge.
@@ 3 Problem Description
Consider a dynamical system described by a multivariate ordinary differential equation (ODE).
$\dot{\mathrm{\mathbf{x}}}(t)=\frac{d\mathrm{\mathbf{x}}(t)}{dt}=\mathbf{f}(\mathrm{\mathbf{x}}(t))$ (1)
where $\mathrm{\mathbf{x}}(t)\in\mathbb{R}^{D}$ denotes the state of a $D$-dimensional system at time $t$ and $\dot{\mathrm{\mathbf{x}}}(t)\in\mathbb{R}^{D}$ is its first order time derivative. The $\mathbf{f}(\mathrm{\mathbf{x}}(t))$ specifies the vector-valued time derivative function. The state of a dynamical system at time $t$ can be obtained by integrating the ODE as:
$\mathrm{\mathbf{x}}(t)=\mathrm{\mathbf{x}}_{0}+\int_{0}^{t}\mathbf{f}(\mathbf{x}(\tau))d\tau$ (2)
The ODE is integrated forward in time starting from the initial condition $\mathbf{x}(0)=\mathbf{x}_{0}$. We assume that the vector field $\mathbf{f}$ is unknown, but it can be learned based on observed data. The goal of a predictive model is to predict $\mathbf{x}(t)$ as accurately as possible, for as long as possible, that is, to minimize the distance between the prediction $\hat{\mathbf{x}}(t)$, and ground-truth trajectory, $\mathbf{x}(t)$. Additionally, in this work, for $D$ predictions of each variable in $\mathrm{\mathbf{x}}(t)$, we wish to create a set of models that correct such predictions, $M_{i}(t)$ for $i\in\{1,\dots,D\}$.
The following exposition describes the data generation process and nomenclature for Predictors, Correctors, trajectory types, errors, and so forth. The method is general and can be used for any $D$-dimensional problem as in equation (1). For ease of exposition and without loss of generality, we select a two-dimensional ODE to describe the inner workings and preliminaries. The Lotka-Volterra (LV) system is used here and in section 4 for illustrative purposes; subsequent sections show experimental results for a variety of systems. LV dynamics can be written as:
@@
The state of this system can be defined as: $\mathbf{\tilde{s}}=(x,y)$. We assume a set of ground-truth $N$ trajectories of $T$ timesteps $I_{N}=\{\tau_{n}\}_{n=1}^{N}$ where $\tau_{n}=\{(\mathbf{s}_{t}^{n},\mathbf{s}_{t+1}^{n})_{t}\}_{t=1}^{T}$ and $\mathbf{s}_{t}^{n}=(x_{t}^{n},y_{t}^{n})$. This dataset can be generated by solving the equations with a numerical ODE solver such as $\mathtt{solve\_ivp}$ from $\mathtt{scipy}$ with random initial conditions. To simulate measurement noise, we perturb each state post-integration with additive Gaussian noise, i.e., $\mathbf{s}=\mathbf{\tilde{s}}+\mathbf{\epsilon}$, where $\mathbf{\epsilon}\sim\mathcal{N}(0,\sigma*\Sigma)$ and $\Sigma=\text{diag}(\sigma_{x},\sigma_{y})$. The $\sigma$ is the scaling factor. We name $I_{N}$ an Integrated dataset. We assume that there is a Predictor $B$ (a learned representation of vector-valued function $\mathbf{f}$), trained on $I_{N}$, that predicts $\mathbf{\bar{s}}_{t+1}$  [^1] The overhead bar over a variable shows that it belongs to Autoregressive dataset. given $\mathbf{\bar{s}}_{t}$, autoregressively generating a predicted set of $N$ trajectories starting from the same initial conditions. Precisely, $A_{N}=\{\bar{\tau}_{n}\}_{n=1}^{N}$ where $\bar{\tau}_{n}=\{(\mathbf{\bar{s}}_{t}^{n},\mathbf{\bar{s}}_{t+1}^{n})_{t}\}_{t=1}^{T}$ and $\mathbf{\bar{s}}_{t}^{n}=(\bar{x}_{t}^{n},\bar{y}_{t}^{n})$. We call $A_{N}$ an Autoregressive dataset. The error ($\mathbf{e}_{t}^{n}$) is the difference between $t^{th}$ timestep of $n^{th}$ trajectory from Integrated and Autoregressive, that is, $\mathbf{e}_{t}^{n}=(e_{x_{t}^{n}},e_{y_{t}^{n}})$ where $\mathbf{e}_{t}^{n}=\mathbf{s}_{t}^{n}-\bar{\mathbf{s}}_{t}^{n}$. Therefore, the set of errors of $N$ trajectories is $F_{N}=\{\{\mathbf{e}_{t}^{n}\}_{t=1}^{T}\}_{n=1}^{N}$.
@@ 4 HopCast
The proposed methodology, HopCast, consists of a Predictor-Corrector mechanism shown in Fig. 1. At any timestep ($t$) during autoregression, the Predictor $B$ produces a point forecast of the next state of the system $(\bar{x}_{t+1},\bar{y}_{t+1})$ given previous state $(\bar{x}_{t},\bar{y}_{t})$. The Corrector retrieves context-dependent errors ($E_{\bar{x}}$ and $E_{\bar{y}}$) to correct the forecast and quantify uncertainty. In particular, the Corrector consists of correction models ($M_{x}$ and $M_{y}$), one for each output of the Predictor $B$. Each correction model ($M$) consists of an Encoder ($m$) and an MHN (ramsauer2020hopfield; auer2023conformal). The Encoder ($m$) is a fully connected feedforward model. Architectural details are included in Appendix E.1. We frame correction as a pattern retrieval task (ramsauer2020hopfield). Precisely, $M_{x}$ and $M_{y}$ take in the context state $(\bar{x}_{1},\bar{y}_{1},\bar{x}_{t},\bar{y}_{t},t)$ as a pattern and outputs $E_{\bar{x}}$ and $E_{\bar{y}}$, respectively. $E_{\bar{x}}$ and $E_{\bar{y}}$ contain errors associated with similar context states from the past. The context state is the input of Predictor $B$ $(\bar{x}_{t},\bar{y}_{t})$ augmented with the initial condition $(\bar{x}_{1},\bar{y_{1}})$ and time step ($t$). The context state captures both local and trajectory-level structure, enabling the Corrector to retrieve errors specific to the current dynamical regime. The $E_{\bar{x}}$ & $E_{\bar{y}}$ are used to derive prediction intervals.
@@
The remainder of this section consists of three subsections. Section 4.1 discusses the shuffling of $I_{N}$, $A_{N}$, & $F_{N}$ datasets to train correction models for pattern retrieval tasks. Section 4.2 then discusses the training of the correction model ($M_{x}$), and section 4.3 describes the inference with $M_{x}$ where we generate the set of errors ($E_{\bar{x}}$) given the context state.
@@ 4.1 Shuffled Data
The MHN in HopCast is trained as a pattern–retrieval model rather than a sequence model. To this end, we break every trajectory into independent context–error pairs and shuffle them, so the model learns to retrieve errors based on the similarity of context, rather than the temporal ordering of the data. We explain that below.
Shuffled Queries ($\mathcal{S}_{Q}$). Before training the correction model ($M_{x}$), the states of the system from Autoregressive ($A_{N}$) dataset are augmented with context information, that is, initial condition and time step. Appendix E.2 emphasizes the importance of adding initial condition as a context. Specifically, the $t^{th}$ timestep of $n^{th}$ trajectory $\bar{\tau}_{n}$ from $A_{N}$, i.e., $(\bar{x}_{t}^{n},\bar{y}_{t}^{n})$, will be $\bar{\mathbf{c}}_{t}^{n}=(\bar{x}_{1}^{n},\bar{y}_{1}^{n},\bar{x}_{t}^{n},\bar{y}_{t}^{n},t)$ after adding context information. The tuple $(\bar{x}_{1}^{n},\bar{y}_{1}^{n})$ is the initial condition of trajectory $\bar{\tau}_{n}$. The $\bar{\mathbf{c}}_{t}^{n}$ denotes the context state associated with the $t^{th}$ timestep of $n^{th}$ trajectory. The set of all context states constructed from dataset $A_{N}$ is $\{\{\bar{\mathbf{c}}_{t}^{n}\}_{t=1}^{T}\}_{n=1}^{N}$. This nested set of context states is flattened and shuffled randomly to form a new set $\mathcal{S}_{Q}=\text{Shuffle}\left(\bigcup_{n=1}^{N}\{\bar{\mathbf{c}}_{t}^{n}\}_{t=1}^{T}\right)$. Appendix E.7 highlights the importance of shuffling the data.
Shuffled Keys ($\mathcal{S}_{K}$). Likewise, we construct the context states $\{\{\mathbf{c}_{t}^{n}\}_{t=1}^{T}\}_{n=1}^{N}$ for Integrated ($I_{N}$) dataset as well, where $\mathbf{c}_{t}^{n}=(x_{1}^{n},y_{1}^{n},x_{t}^{n},y_{t}^{n},t)$. Similar to $\mathcal{S}_{Q}$, we flatten the nested set $\{\{\mathbf{c}_{t}^{n}\}_{t=1}^{T}\}_{n=1}^{N}$ and randomly shuffle it to form $\mathcal{S}_{K}=\text{Shuffle}\left(\bigcup_{n=1}^{N}\{\mathbf{c}_{t}^{n}\}_{t=1}^{T}\right)$.
@@
Once we have $\mathcal{S}_{Q}$, $\mathcal{S}_{K}$, $\mathcal{S}_{V_{x}}$, $\mathcal{S}_{V_{y}}$, they are split into train/test with an 80/20 split. To train $M_{x}$, we need dataset $\mathcal{D}_{x}=\{(\mathbf{Q},\mathbf{K},\mathbf{V}_{x})_{i}\}$ comprising triplets $(\mathbf{Q},\mathbf{K},\mathbf{V}_{x})$. Each triplet of $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}_{x}$ are constructed by sampling $\mathtt{S_{L}}$ (Sequence Length) number of elements from sets $\mathcal{S}_{Q}^{\text{Train}}$, $\mathcal{S}_{K}^{\text{Train}}$, and $\mathcal{S}_{V_{x}}^{\text{Train}}$, respectively. For instance, one triplet with $\mathtt{S_{L}}=5$ is: $\mathbf{Q}=(\bar{\mathbf{c}}_{4}^{3},\bar{\mathbf{c}}_{5}^{1},\bar{\mathbf{c}}_{7}^{5},\bar{\mathbf{c}}_{2}^{3},\bar{\mathbf{c}}_{1}^{7})^{\top}$, $\mathbf{K}=(\mathbf{c}_{4}^{3},\mathbf{c}_{5}^{1},\mathbf{c}_{7}^{5},\mathbf{c}_{2}^{3},\mathbf{c}_{1}^{7})^{\top}$, and $\mathbf{V}_{x}=(e_{x_{4}^{3}},e_{x_{5}^{1}},e_{x_{7}^{5}},e_{x_{2}^{3}},e_{x_{1}^{7}})^{\top}$. The error $e_{x_{4}^{3}}=x_{4}^{3}-\bar{x}_{4}^{3}$ denotes the error of Predictor $B$ at $4^{th}$ timestep of $3^{rd}$ trajectory. Likewise, we construct $\mathcal{D}_{y}=\{(\mathbf{Q},\mathbf{K},\mathbf{V}_{y})_{i}\}$ using $\mathcal{S}_{Q}^{\text{Train}}$, $\mathcal{S}_{K}^{\text{Train}}$ and $\mathcal{S}_{V_{y}}^{\text{Train}}$. The 20% split with inference data is used for final evaluation.